OpenStudy (anonymous):

i need help please!!!!!!!!!!! find an equation of the line having the given slope and containing the given points slope:2; through (9,5)

pooja195 (pooja195):

y-y1=m(x-x1)

pooja195 (pooja195):

m=slope (x1,y1)

OpenStudy (anonymous):

slope is 2 tho

OpenStudy (anonymous):

so wouldnt it be y=2x+

pooja195 (pooja195):

right m is just the vairble for it :P \[\huge~y-5=2(x-9)\]

pooja195 (pooja195):

2(x−9) distribute

OpenStudy (anonymous):

yes

pooja195 (pooja195):

What did you get?

OpenStudy (anonymous):

then add 5?

pooja195 (pooja195):

First distribute so i know you're doing it right .

OpenStudy (anonymous):

lol 2x-18

pooja195 (pooja195):

\[\huge~y-5=2x-18\] good now add 5

pooja195 (pooja195):

\[\huge~-18+5=?\]

OpenStudy (anonymous):

y=2x-13

pooja195 (pooja195):

and you have your answer! :) good job :)

OpenStudy (anonymous):

thanks!